In a shipment of 78 vials, only 18 do not have airline cracks. If you randomly select 3 vials from the shipment, what is the probability that none of the 3 vials have airlines? what is the probability that all 3 selected have airlines?
Answer by stanbon(75887)

In a shipment of 78 vials, only 18 do not have hairline cracks. If you randomly select 3 vials from the shipment,
P(have crack) = 18/78
P(no crack) = 60/78
--------------------------------------
what is the probability that none of the 3 vials have hairlines?
Ans: (60/78)^3 = 0.4552
------------------------------------
what is the probability that all 3 selected have hairlines?
Ans: (18/78)^3 = 0.0123
================================
Cheers,
Stan H.